OpenStudy (anonymous):

how do i solve 2/7 = x/-3 using the multiplication principle

OpenStudy (tkhunny):

The multiplication principle is "multiplication"? \(\dfrac{2}{7} = \dfrac{x}{-3}\) Multiply by 7 \(7\cdot\dfrac{2}{7} = 7\cdot\dfrac{x}{-3}\) Simplify \(2 = 7\cdot\dfrac{x}{-3}\) Multiply by -3 \(-3*2 = -3*7\cdot\dfrac{x}{-3}\) Simplify \(-6 = 7\cdot x\) Yeah! Multiplication still works!!! We don't need another principle for that.